Find the length of the portion of the parabola x =3t^2,y=6t cut off by the line 3x+y-3=0
1
Expert's answer
2020-03-10T12:52:50-0400

The length of the curve is calculated by the formula:


"l=\int^{x_2}_{x_1}\sqrt{1+(f'(x))^2}dx"

wrere "x_1" , "x_2" - curve borders.


Bring the parabola to the form "x=f(y)":

"y=6t\rarr t=\frac{y}{6}" ;

"x=3t^2=\frac{y^2}{12}" .


Find the boundaries of the curve as the intersection of two functions.

"\begin{cases}
 3x+y-3=0 \\
 x=\frac{y^2}{12}
\end{cases}\rarr \frac{y^2}{4}+y-3=0 \rarr y_1=-6, y_2=2".


Calculate the derivative of the function "f(y)" :

"f'(y)=(\frac{y^2}{12})'=\frac{y}{6}\rarr (f'(y))^2=\frac{y^2}{36}" .


We calculate the resulting integral


"\int^{2}_{-6}\sqrt{1+\frac{y^2}{36}}dy=\frac{1}{6}\int^{2}_{-6}\sqrt{36+y^2}dy"


Use hyperbolic substitution

"y=6\sinh{u} , \sqrt{36+y^2}=6\cosh{u},"

"dy=6\cosh{u}\space du, u=arcsinh({\frac{y}{6}}),"

"u_1=arcsin({\frac{y_1}{6}})=-arcsinh(1),""u_2=arcsin({\frac{y_2}{6}})=arcsinh(\frac{1}{3})."

We have

"\int^{2}_{-6}\sqrt{1+\frac{y^2}{36}}dy=\frac{1}{6}\int_{-arcsin(1)}^{arcsin(\frac{1}{3})}36\cosh^2{u}\space du=6\int_{-arcsin(1)}^{arcsin(\frac{1}{3})}\frac{\cosh{2u}+1}{2}du=6(\frac{1}{4}\sinh{2u}+\frac{u}{2})|^{arcsin(\frac{1}{3})}_{-arcsin(1)}=(3\sinh{u}\space \cosh{u}+3u)|^{arcsin(\frac{1}{3})}_{-arcsin(1)}=\frac{\sqrt{10}}{3}+3\sqrt2+3arcsinh(\frac{1}{3})+3arcsinh(1)≈8.923"